OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 3-inch radius so for this i had 27/3sqrt4 which is probally not right

jhonyy9 (jhonyy9):

area =(base *h)/2 - yes but we know that the radius is equal 2h/3 from what we can get h=3r/2 - so being the length of one side l - in one triangle equilateral h half the base on what is perpendicular ,so in this way we get are for half part of triangle will be ((l/2)*h)/2 = (lh/2)/2 = lh/4 ,hence the complete area will be 2(lh/4)=2lh/4 =lh/2 but we know that h=3r/2 , so hence area =l(3r/2) = =3lr/2 ---- A=3lr/2 - if A=3lr/2 and r=3 inch hence A=9l/2